# Chain-of-Thought Prompting for Speech Translation

### v1

@@ Abstract
Large language models (LLMs) have demonstrated remarkable advancements in language understanding and generation. Building on the success of text-based LLMs, recent research has adapted these models to use speech embeddings for prompting, resulting in Speech-LLM models that exhibit strong performance in automatic speech recognition (ASR) and automatic speech translation (AST). In this work, we propose a novel approach to leverage ASR transcripts as prompts for AST in a Speech-LLM built on an encoder-decoder text LLM. The Speech-LLM model consists of a speech encoder and an encoder-decoder structure Megatron-T5. By first decoding speech to generate ASR transcripts and subsequently using these transcripts along with encoded speech for prompting, we guide the speech translation in a two-step process like chain-of-thought (CoT) prompting. Low-rank adaptation (LoRA) is used for the T5 LLM for model adaptation and shows superior performance to full model fine-tuning. Experimental results show that the proposed CoT prompting significantly improves AST performance, achieving an average increase of 2.4 BLEU points across 6 EnX or XEn AST tasks compared to speech prompting alone. Additionally, compared to a related CoT prediction method that predicts a concatenated sequence of ASR and AST transcripts, our method performs better by an average of 2 BLEU points.
Index Terms— Chain of thought, prompting, ASR, AST, LLM
1 Introduction
@@
Past work in various speech tasks also shows the benefits of a multi-step prediction. For example, deliberation models [21, 22, 23, 24] first predict the first-pass hypothesis and then use that to assist a more sophisticated second-pass task. In the Speech-LLM framework, a joint audio and speech understanding model [25] uses Whisper [26] to generate spoken text and use that to prompt the LLaMA LLM [5] for a range of audio tasks. Without using speech in prompting, [27] develops a generic multi-task correction LLM takes outputs from various models and generates refined results. Recently, in [28], a CoT prediction method for speech translation is proposed to predict a concatenated sequence of ASR and AST transcripts by prompting a decoder-only GPT. However, when using an encoder-decoder LLM architecture such as T5, it is unclear what is the best place to inject the ASR hypothesis, i.e., decoder outputs (as in [28]) or T5 inputs, and is worth researching.
@@ 2 Model Description
The novelty of the proposed model is discussed by comparing to related systems below. Compared to the baseline model (same as Fig. 1 but without ASR text hypothesis in prompts), our model performs translation by first predicting ASR hypothesis and then use that as input in addition to speech embedding for translation. The ASR prediction acts as the first step in a two-step translation process, similar to the CoT prompting in machine translation [20] where the model is given multiple steps of instructions to guide a task. The use of both speech and ASR predictions in the second-step prediction is similar to deliberation models [22, 23], and here we further capitalize on the power of LLM by prompting. Compared to the CoT prediction method in [28], our model uses an encoder-decoder structure T5 LLM instead of a decoder-only GPT in [28]. When using an encoder-decoder LLM architecture such as T5, it is unclear what is the best place to inject the ASR hypothesis (i.e., decoder outputs or T5 inputs) and is worth investigating. In addition, we present CoT AST results in multiple language pairs instead of English-Chinese translations in [28]. The effectiveness of injecting ASR hypotheses as T5 inputs is presented in Sect. 4.5 by comparing to the CoT prediction method.
@@
3 Experiment Details
3.1 Data
As shown in Table 1, we use a subset of Canary AST training data [31] to generate CoT training data for experiment efficiency. We first generate the ASR hypothesis and then append the hypothesis to the fixed model prompt as T5 input. The combined textual prompt looks like:
@@ 3.1 Data
Here, and are source and target language names, respectively, and represents the estimated ASR text hypothesis for the utterance. The punctuation and capitalization prompts are applied depending on the target text. We have generated ASR hypotheses as described in Sect. 2 for 3 source languages in XEn, and for English in EnX translations. Our AST target labels are generated synthetically by NVIDIA Megatron NMT models [36, 37]. In inference, we evaluate AST model performance using the FLEURS [38] test sets. We first generate ASR hypothesis in the same way as in training and then inject them in AST prompt for inference.
| Lang. Pair | Training Data (hr) |
| DeEn | 2,752 |
| FrEn | 1,795 |
| EsEn | 1,397 |
| EnDe | 1,640 |
| EnFr | 1,640 |
| EnEs | 1,640 |
3.2 Modeling Details
We implement the model with PyTorch using NeMo Tookit [39], and the model is trained on 32 A100 (80G) GPUs with a batch duration of 180 sec per GPU. The speech encoder is initialized from the 20-layer Canary-1B encoder [31], and the LLM is initialized from the 1.2B Megatron-T5 NMT model [35]. The T5 LLM consists of a 12-layer encoder and a 24-layer decoder, with 20 attention heads and a hidden dimension of 1280, and the feedforward layer has a dimension of 5120. Relative positional embedding is used. We use a 64k SentencePiece tokenizer for all languages. RMSNorm [40] is used for normalization for all layers of the T5 model. We use fused Adam, and an inverse Square Root Annealing learning rate (LR) schedule for optimization. The LR schedule starts with an initial learning rate of 4e-4. Gradient clipping is applied at a threshold of 1.0 to stabilize training. Warmup steps are configured to 0.8% of 3.6M maximum steps.
@@ 3.2 Modeling Details
4 Results
In this section, we first present ablation studies for model design, and then compare the best performing model to related work.
4.1 Chain-of-Thought (CoT) prompting
We first evaluate the performance of the CoT prompting model, i.e., injecting the ASR hypotheses to the LLM prompt. As shown in Table 2, the CoT prompt (E1) benefits translation for all language pairs and achieved an average improvement of 1.5 BLEU score compared to the baseline (from 31.1 32.6). The baseline model is trained in the same way as the CoT prompt model, except removing ASR hypotheses and using the prompt such as: “Translate the spoken content to written text, with punctuations and capitalizations.” We call the baseline model SALM-T5 since the prompt concatenation follows the same way as the SALM [10]. Our improvement is for both EnX and XEn translations. In this experiment, both training and inference have used estimated ASR hypotheses. The ASR hypotheses are generated and then appended to the AST prompt as described in Sect. 3.1.
@@ 4.1 Chain-of-Thought (CoT) prompting
4.2 Prompt with ground truth ASR transcripts
To measure the impact of estimated ASR hypothesis quality on CoT prompting, we have tried using ground truth ASR labels in prompts. This is to evaluate whether there is potential improvement by using a better ASR model. As indicated in Table 3, we have obtained an average of 2.7 BLEU score improvement (32.6 35.3) for all languages by using the ground truth ASR labels, compared to E1 which uses estimated ASR hypotheses. It means that better quality ASR prediction does benefit translation. On the other hand, it will be interesting to see how the model performs with lower quality ASR transcripts as inputs, i.e., maybe ones generated from a lightweight and efficient model as the first pass.
@@ 4.2 Prompt with ground truth ASR transcripts
4.3 CoT prediction
Since we use an encoder-decoder T5 instead of the decoder-only GPT, we investigate what is the best place to inject ASR hypotheses. In addition to injecting in the T5 encoder (Sect. 4.1), we have also tried predicting ASR hypotheses first and then followed by AST output (i.e., [28]). Similar to [28], we use the following prompt: “Q: Transcribe the spoken content to written {source_lang} text, then translate this to {target_lang} text, with punctuations and capitalizations.” We have experimented two setups in this experiment: 1) ASR ground truth text is used in the target sequence for prediction (same as [28]), or 2) estimated ASR hypotheses are used target sequence to create a more matched condition as in inference (second column in Table 4). When using ground truth ASR hypotheses, the model predicts the concatenated ASR and AST hypotheses in a single sequence. We use a special separator token to concatenate the two labels, and the loss is calculated for the whole sequence. When using the predicted ASR hypotheses, we mask the loss over the ASR part and only use the loss from the AST prediction. In either scenario (in Table 4), we have not observed better performance of the CoT prediction method compared to the CoT prompting method.
@@ 4.4 Low-rank adaptation (LoRA) performance
We have experimented adding LoRA to the Megatron-T5 model and achieved significant improvements across all languages (Table 5). Since the Megatron T5 has an encoder-decoder structure, LoRA adapters have been added to both the encoder and decoder. We use an adapter dimension of 128 for all the 12 self-attention layers in the encoder. For the decoder, we have added LoRA adapters to both self-attention and cross attention layers for every decoder layer. For self-attention layers, we use the same adapter dimension of 128, and for cross-attention, we use an adapter dimension of 32 for queries and 64 for keys and values. In total, this adds 8M and 47M parameters to the encoder and decoder, respectively. As shown in Table 5, the LoRA adapter significantly improves the model performance by improving the BLEU score from 32.6 33.5. The improvement ranges from 0.3 to 1.8 BLEU. LoRA seems to maintain the translation ability of the original text LLM, which may be beneficial in our scenario with text ASR transcripts in the prompt.
4.5 Comparisons
@@
In Table 6, we compare our CoT prompting models (E1 and E2) to several models including the SALM-T5 baseline (B1), a CoT prediction model (B2), and SeamlessM4T models (B3 and B4). Note that we have implemented the CoT prediction [28] based on the encoder-decoder T5 LLM instead of the decoder-only GPT in [28]. For SeamlessM4T [41], we have used the official checkpoints to rerun the model. We evaluate the models by using the FLEURS dataset [38] for translation of 6 language pairs (3 XEn and 3 EnX).
@@
We first compare the SALM-T5 baseline (B1) and the CoT prompting (E1), and the performance difference is only due to adding ASR hypotheses to the T5 prompts. We see in Table 6 that the latter performs 1.5 BLEU score better on average. Then, by adding LoRA adapters, our CoT prompting model performs an additional 0.9 BLEU better, achieving an average of 2.4 BLEU better compared to the baseline (B1). We can see the improvement is uniform for all language pairs. Secondly, we compare to the CoT prediction method in [28]. Compared to the baseline B1, the results show that CoT prediction [28] improves the baseline by around 0.4 BLEU. It is effective for most language pairs but worse in DeEn translations. Comparing CoT prediction (B2) to the proposed CoT prompting method (E1), the injection of ASR hypotheses in prompts improves the BLEU by 0.9 on average (31.5 32.6). This improvement shows that injecting the ASR transcripts as T5 inputs is more effective than decoder outputs. Note that we have used the same ASR hypotheses in both models. Lastly, to put model performance into perspective, we compare our best performing model (E2, with LoRA adapters) to the SeamlessM4T medium and large models [41]. Our LoRA model performs 4.7 and 1.3 BLEU better than the medium and large SeamlessM4T models, respectively. The SeamlessM4T medium and large models have sizes of 1.2B and 2.3B, respectively, while our model has a total size of 1.8B. However, we also note that the proposed model only performs speech translation but SeamlessM4T is capable of a range of speech and text tasks.
@@
To measure the effect of using both speech and ASR text transcription for prompting, we further compare our CoT prompting + LoRA model (E2) to a cascade system. In the cascade system, the ASR hypotheses are first generated and then a machine translation model is used to translate the hypotheses to target texts. We have used the same ASR hypotheses as those for CoT prompting in E2 for fair comparison. For machine translation, we have used two separate Megatron 500M models for En-to-any [37] and any-to-En [36] translations, totaling 1B model parameters. We have tried using a 1B any-to-any MT model for translation but achieved worse results. In Table 7, we can see that our system performs similarly in 3 XEn language pairs, and 0.5-2.5 BLEU score better for EnX translations. The better performance of our model in translating English to other languages is probably due to initializing from the Canary encoder, which has more training data in English than other languages. On average, our CoT prompting + LoRA model performs 0.7 BLEU better than the cascade system.
5 Conclusions
@@
We have investigated chain-of-thought (CoT) prompting for a Speech-LLM built on an encoder-decoder architecture LLM. The ASR transcripts are first estimated and then injected in the model prompts for the second-step of translation. Our best performing model, incorporating LoRA, achieved an average BLEU score improvement of 2.4 points compared to the SALM-T5 baseline. Compared to a CoT prediction model similar to [28], our method performs 2 BLEU score better across all language pairs, demonstrating the effectiveness of utilizing ASR transcripts as inputs for a encoder-decoder LLM. Although our model is trained using limited amounts of data, it is competitive in speech translation compared to related models such as SeamlessM4T. The effectiveness of using both speech and ASR text in prompting is demonstrated by a gain of up to 2.5 BLEU over a traditional cascade system. We will release the code and checkpoints to promote the next-generation SpeechLLM design.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable advancements in language understanding and generation. Building on the success of text-based LLMs, recent research has adapted these models to use speech embeddings for prompting, resulting in Speech-LLM models that exhibit strong performance in automatic speech recognition (ASR) and automatic speech translation (AST). In this work, we propose a novel approach to leverage ASR transcripts as prompts for AST in a Speech-LLM built on an encoder-decoder text LLM. The Speech-LLM model consists of a speech encoder and an encoder-decoder structure Megatron-T5. By first decoding speech to generate ASR transcripts and subsequently using these transcripts along with encoded speech for prompting, we guide the speech translation in a two-step process like chain-of-thought (CoT) prompting. Low-rank adaptation (LoRA) is used for the T5 LLM for model adaptation and shows superior performance to full model fine-tuning. Experimental results show that the proposed CoT prompting significantly improves AST performance, achieving an average increase of 2.4 BLEU points across 6 EnX or XEn AST tasks compared to speech prompting alone. Additionally, compared to a related CoT prediction method that predicts a concatenated sequence of ASR and AST transcripts, our method performs better by an average of 2 BLEU points. Our code is open-sourced on GitHub.111https://github.com/kevinhu-nv/NeMo_CoT
I Introduction
@@
Past work in various speech tasks also shows the benefits of a multi-step prediction. For example, deliberation models [21, 22, 23, 24] first predict the first-pass hypothesis and then use that to assist a more sophisticated second-pass task. In the Speech-LLM framework, a joint audio and speech understanding model [25] uses Whisper [26] to generate spoken text and use that to prompt the LLaMA LLM [5] for a range of audio tasks. Without using speech in prompting, [27] develops a generic multi-task correction LLM takes outputs from various models and generates refined results. Recently, in [28], a CoT prediction method for speech translation is proposed to predict a concatenated ASR and AST transcript sequence by prompting a decoder-only GPT. However, when using an encoder-decoder LLM such as T5, it is unclear what is the best place to inject the ASR hypothesis, i.e., decoder outputs (as in [28]) or T5 inputs, and is worth researching (note [29], [30] are developed concurrently with our work).
@@ II Model Description
The novelty of the proposed model is discussed by comparing to related systems below. Compared to the baseline model (same as Fig. 1 but without ASR text hypothesis in prompts), our model performs translation by first predicting ASR hypothesis and then use that as input in addition to speech embedding for translation. The ASR prediction acts as the first step in a two-step translation process, similar to the CoT prompting in machine translation [20] where the model is given multiple steps of instructions to guide a task. The use of both speech and ASR predictions in the second-step prediction is similar to deliberation models [22, 23], and here we further capitalize on the power of LLM by prompting. Compared to the CoT prediction method in [28], our model uses an encoder-decoder structure T5 LLM instead of a decoder-only GPT in [28]. When using an encoder-decoder LLM architecture such as T5, it is unclear what is the best place to inject the ASR hypothesis (i.e., decoder outputs or T5 inputs) and is worth investigating. In addition, we present CoT AST results in multiple language pairs instead of English-Chinese translations in [28]. The effectiveness of injecting ASR hypotheses as T5 inputs is presented in Sect. IV-E by comparing to the CoT prediction method.
@@
III Experiment Details
III-A Data
We use a subset of Canary AST training data [33] to generate CoT training data for experiment efficiency (XEn with 2,752, 1,795, and 1,397 hours for De, Fr, and Es, respectively, and EnX with 1,640 hours for each language of De, Fr, and Es). We first generate the ASR hypothesis and then append the hypothesis to the fixed model prompt as T5 input. The combined textual prompt looks like:
@@ III-A Data
Here, and are source and target language names, respectively, and represents the estimated ASR text hypothesis for the utterance. The punctuation and capitalization prompts are applied depending on the target text. We have generated ASR hypotheses as described in Sect. II for 3 source languages in XEn, and for English in EnX translations. Our AST target labels are generated synthetically by NVIDIA Megatron NMT models [38, 39]. In inference, we evaluate AST model performance using the FLEURS [40] test sets. We first generate ASR hypothesis in the same way as in training and then inject them in AST prompt for inference.
III-B Modeling Details
We implement the model with PyTorch using NeMo Toolkit [41], and the model is trained on 32 A100 (80G) GPUs with a batch duration of 180 sec per GPU. The speech encoder is initialized from the 20-layer Canary-1B encoder [33], and the LLM is initialized from the 1.2B Megatron-T5 NMT model [37]. The T5 LLM consists of a 12-layer encoder and a 24-layer decoder, with 20 attention heads and a hidden dimension of 1280, and the feedforward layer has a dimension of 5120. Relative positional embedding is used. We use a 64k SentencePiece tokenizer for all languages. RMSNorm [42] is used for normalization for all layers of the T5 model. We use fused Adam, and an inverse Square Root Annealing learning rate (LR) schedule for optimization. The LR schedule starts with an initial learning rate of 4e-4. Gradient clipping is applied at a threshold of 1.0 to stabilize training. Warmup steps are configured to 0.8% of 3.6M maximum steps.
@@ III-B Modeling Details
IV Results
IV-A Chain-of-Thought (CoT) prompting
@@
We first evaluate the performance of the CoT prompting model, i.e., injecting the ASR hypotheses to the LLM prompt. As shown in Table I, the CoT prompt (E1) benefits translation for all language pairs and achieved an average improvement of 1.5 BLEU score compared to the baseline (from 31.1 32.6). The baseline model is trained in the same way as the CoT prompt model, except removing ASR hypotheses and using the prompt such as: “Translate the spoken content to written text, with punctuations and capitalizations.” We call the baseline model SALM-T5 since the prompt concatenation follows the same way as the SALM [10]. Our improvement is for both EnX and XEn translations. In this experiment, both training and inference have used estimated ASR hypotheses. The ASR hypotheses are generated and then appended to the AST prompt as described in Sect. III-A.
@@
IV-B Prompt with ground truth ASR transcripts
To measure the impact of estimated ASR hypothesis quality on CoT prompting, we have tried using ground truth ASR labels in prompts. This is to evaluate whether there is potential improvement by using a better ASR model. As indicated in Table II, we have obtained an average of 2.7 BLEU score improvement (32.6 35.3) for all languages by using the ground truth ASR labels, compared to E1 which uses estimated ASR hypotheses. It means that better quality ASR prediction does benefit translation. On the other hand, it will be interesting to see how the model performs with lower quality ASR transcripts as inputs, i.e., maybe ones generated from a lightweight and efficient model as the first pass.
@@ IV-B Prompt with ground truth ASR transcripts
IV-C CoT prediction
Since we use an encoder-decoder T5 instead of the decoder-only GPT, we investigate what is the best place to inject ASR hypotheses. In addition to injecting in the T5 encoder (Sect. IV-A), we have also tried predicting ASR hypotheses first and then followed by AST output (i.e., [28]). Similar to [28], we use the following prompt: “Q: Transcribe the spoken content to written {source_lang} text, then translate this to {target_lang} text, with punctuations and capitalizations.” We have experimented two setups in this experiment: 1) ASR ground truth text is used in the target sequence for prediction (same as [28]), or 2) estimated ASR hypotheses are used target sequence to create a more matched condition as in inference (second column in Table III). When using ground truth ASR hypotheses, the model predicts the concatenated ASR and AST hypotheses in a single sequence. We use a special separator token to concatenate the two labels, and the loss is calculated for the whole sequence. When using the predicted ASR hypotheses, we mask the loss over the ASR part and only use the loss from the AST prediction. In either scenario (in Table III), we have not observed better performance of the CoT prediction method compared to the CoT prompting method.
IV-D Low-rank adaptation (LoRA) performance
@@
We have experimented adding LoRA to the Megatron-T5 model and achieved significant improvements across all languages (Table IV). Since the Megatron T5 has an encoder-decoder structure, LoRA adapters have been added to both the encoder and decoder. We use an adapter dimension of 128 for all the 12 self-attention layers in the encoder. For the decoder, we have added LoRA adapters to both self-attention and cross attention layers for every decoder layer. For self-attention layers, we use the same adapter dimension of 128, and for cross-attention, we use an adapter dimension of 32 for queries and 64 for keys and values. In total, this adds 8M and 47M parameters to the encoder and decoder, respectively. As shown in Table IV, the LoRA adapter significantly improves the model performance by improving the BLEU score from 32.6 33.5. The improvement ranges from 0.3 to 1.8 BLEU. LoRA seems to maintain the translation ability of the original text LLM, which may be beneficial in our scenario with text ASR transcripts in the prompt.
IV-E Comparisons
@@
In Table V, we compare our CoT prompting models (E1 and E2) to several models including the SALM-T5 baseline (B1), a CoT prediction model (B2), and SeamlessM4T models (B3 and B4). Note that we have implemented the CoT prediction [28] based on the encoder-decoder T5 LLM instead of the decoder-only GPT in [28]. For SeamlessM4T [43], we have used the official checkpoints to rerun the model. We evaluate the models by using the FLEURS dataset [40] for translation of 6 language pairs (3 XEn and 3 EnX).
@@
We first compare the SALM-T5 baseline (B1) and the CoT prompting (E1), and the performance difference is only due to adding ASR hypotheses to the T5 prompts. We see in Table V that the latter performs 1.5 BLEU score better on average. Then, by adding LoRA adapters, our CoT prompting model performs an additional 0.9 BLEU better, achieving an average of 2.4 BLEU better compared to the baseline (B1). We can see the improvement is uniform for all language pairs. Secondly, we compare to the CoT prediction method in [28]. Compared to the baseline B1, the results show that CoT prediction [28] improves the baseline by around 0.4 BLEU. It is effective for most language pairs but worse in DeEn translations. Comparing CoT prediction (B2) to the proposed CoT prompting method (E1), the injection of ASR hypotheses in prompts improves the BLEU by 0.9 on average (31.5 32.6). This improvement shows that injecting the ASR transcripts as T5 inputs is more effective than decoder outputs. Note that we have used the same ASR hypotheses in both models. Lastly, to put model performance into perspective, we compare our best performing model (E2, with LoRA adapters) to the SeamlessM4T medium and large models [43]. Our LoRA model performs 4.7 and 1.3 BLEU better than the medium and large SeamlessM4T models, respectively. The SeamlessM4T medium and large models have sizes of 1.2B and 2.3B, respectively, while our model has a total size of 1.8B. However, we also note that the proposed model only performs speech translation but SeamlessM4T is capable of a range of speech and text tasks.
@@
To measure the effect of using both speech and ASR text transcription for prompting, we further compare our LoRA model (E2) to a cascade system. In the cascade system, the ASR hypotheses are first generated and then a machine translation model is used for text translation. We have used the same ASR hypotheses as those for CoT prompting in E2 for fair comparison. For machine translation, we have used two separate Megatron 500M models for En-to-any [39] and any-to-En [38] translations, totaling 1B model parameters. We have tried using a 1B any-to-any MT model for translation but achieved worse results. In Table VI, we can see that our system performs similarly in 3 XEn language pairs, and 0.5-2.5 BLEU score better for EnX translations. The better performance of our model in translating English to other languages is probably due to initializing from the Canary encoder.
V Conclusions
@@
We have investigated chain-of-thought (CoT) prompting for a SpeechLLM built on an encoder-decoder architecture. Our best performing model achieved an average BLEU score improvement of 2.4 points compared to the SALM-T5 baseline. Compared to a CoT prediction model similar to [28], our method performs 2 BLEU score better across all language pairs. The effectiveness of using both speech and ASR text in prompting is demonstrated by a gain of up to 2.5 BLEU over a traditional cascade system.